### v2

# The Wavefront Set: Bounds for the Langlands parameter

Abstract.
We are interested in the interplay between the Langlands parameter of a smooth irreducible representation of the reductive $p$-adic group and the geometric wavefront set of the representation, an invariant that appears naturally in the local character expansion of Howe [Ho] and Harish-Chandra [HC].
1. Introduction: a conjecture
Let ${\mathsf{k}}$ be a nonarchimedean field with residue field $\mathbb{F}_{q}$. Let $\bar{\mathsf{k}}$ be a separable closure of ${\mathsf{k}}$ and let $K\subset\bar{\mathsf{k}}$ denote the maximal unramified extension. Let $W_{\mathsf{k}}$ be the Weil group with inertia subgroup $I_{\mathsf{k}}$ and norm ||||||~{}||| | | |. Let $W^{\prime}_{\mathsf{k}}$ be the Weil-Deligne group, i.e., $W^{\prime}_{\mathsf{k}}=W_{\mathsf{k}}\ltimes{\mathbb{C}}$, where the action is defined by $w\cdot z=||w||z$, $w\in W_{\mathsf{k}}$, $z\in{\mathbb{C}}$. Let $\mathsf{Fr}$ denote the geometric Frobenius element, which generates $W_{\mathsf{k}}/I_{\mathsf{k}}$.
Let ${\mathbf{G}}$ be a reductive ${\mathsf{k}}$-group that splits over an unramified extension. The pure inner forms of ${\mathbf{G}}$, in the sense of [Vo] are parametrized by the first Galois cohomology group $H^{1}({\mathsf{k}},{\mathbf{G}})$. For every $\omega\in H^{1}({\mathsf{k}},{\mathbf{G}})$, let ${\mathbf{G}}^{\omega}({\mathsf{k}})$ denote the corresponding pure inner form. Let $G^{\vee}$ be the complex dual group to ${\mathbf{G}}$. Let ${}^{L}\!G=G^{\vee}\rtimes W_{\mathsf{k}}$ be the L-group.
1.1. Langlands parameters
We briefly recall the conjectural Langlands parametrization of the irreducible ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representations, in the formulation of [Vo], generalizing the Kazhdan-Lusztig geometric parametrization of representations with Iwahori-fixed vectors. A Langlands parameter is a group homomorphism
$\varphi:W^{\prime}_{\mathsf{k}}\to{}^{L}\!G$ (1.1)
satisfying certain compatibility conditions. Let $A_{\varphi}$ denote the group of components of the centralizer in $G^{\vee}$ of the image of $\varphi$. One expects a one-to-one correspondence
$\sqcup_{\omega\in H^{1}({\mathsf{k}},{\mathbf{G}})}\mathsf{Irr}{\mathbf{G}}^{\omega}({\mathsf{k}})\leftrightarrow G^{\vee}\text{-conjugacy classes of }(\varphi,\phi),$
where $\varphi$ is a Langlands parameter and $\phi$ is an irreducible representation of $A_{\varphi}$. The pairs $(\varphi,\phi)$ are called enhanced Langlands parameters. Let $\Pi(\varphi)$ denote the Langlands packet, i.e., the set of representations parametrized by a fixed ($G^{\vee}$-orbit of) $\varphi$. We need the geometric reformulation of this correspondence.
Fix a group homomorphism $\lambda:I_{\mathsf{k}}\to{}^{L}\!G$ with finite image. The enhanced Langlands parameters for which $\varphi|_{I_{\mathsf{k}}}=\lambda$ admit the following interpretation. Let $H^{\vee}={G^{\vee}}^{\lambda(I_{\mathsf{k}})}$, a possibly disconnected reductive group, and ${\mathfrak{h}}^{\vee}={\mathfrak{g}^{\vee}}^{\lambda(I_{\mathsf{k}})}$, a reductive subalgebra of $\mathfrak{g}^{\vee}$, denote the fixed points.
Let $s=\varphi(\mathsf{Fr})\in{}^{L}\!G$ be a semisimple element. Since $\mathsf{Fr}$ normalizes $I_{\mathsf{k}}$, it follows that $s$ normalizes $H^{\vee}$ (and ${\mathfrak{h}}^{\vee}$). Let ${\mathfrak{h}}^{\vee}_{q}$ denote the $q$-eigenspace of $\mathrm{Ad}(s)$. Let $H^{\vee}(s)$ be the centralizer of $s$ in $H^{\vee}$. Then $H^{\vee}(s)$ acts on ${\mathfrak{h}}^{\vee}_{q}$ with finitely many orbits. There is a one-to-one correspondence between $G^{\vee}$-orbits of enhanced Langlands parameters $(\varphi,\phi)$ such that $\varphi(\mathsf{Fr})=s$ and $\varphi|_{I_{\mathsf{k}}}=\lambda$, and pairs $({{\mathcal{O}}}^{\vee},\mathcal{L}^{\vee})$ where
  • ${{\mathcal{O}}}^{\vee}$ is an $H^{\vee}(s)$-orbit in ${\mathfrak{h}}^{\vee}_{q}$;
  • $\mathcal{L}^{\vee}$ is an irreducible $H^{\vee}(s)$-equivariant local system supported on ${{\mathcal{O}}}^{\vee}$.
Denote by $\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})=\Pi(\varphi)$ the corresponding L-packet, each of whose member $\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ should be parametrized by the equivariant local systems $\mathcal{L}^{\vee}$ on ${{\mathcal{O}}}^{\vee}$. A desideratum is that $\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})$ consists of tempered representations if and only if ${{\mathcal{O}}}^{\vee}$ is dense in ${\mathfrak{h}}^{\vee}_{q}$.
1.2. Aubert-Zelevinsky involution
Let ${\mathsf{AZ}}$ denote the normalized Aubert-Zelevinsky involution [Au, Definition 1.5, Corollary 3.9] on $\mathsf{Irr}{\mathbf{G}}^{\omega}({\mathsf{k}})$ for every form ${\mathbf{G}}^{\omega}({\mathsf{k}})$. If $\varphi_{\pi}$ is the Langlands parameter of $\pi$, then it is expected that $\varphi_{\pi}(W_{\mathsf{k}})$ and $\varphi_{{\mathsf{AZ}}(\pi)}(W_{\mathsf{k}})$ are $G^{\vee}$ conjugate. Moreover, if $\pi=\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$, it is expected that the Langlands parameter of ${\mathsf{AZ}}(\pi)$ is related to that of $\pi$ via the Fourier transform [KS, §3.7] on the appropriate space of perverse sheaves of Langlands parameters, see also [EM, §3]. This is known to hold for all representations with Iwahori-fixed vectors by [EM, Theorem 0.1], and more generally it is expected to hold for unipotent representations in the sense of Lusztig.
More precisely, suppose that $\pi$ has Iwahori-fixed vectors, so $\lambda(I_{\mathsf{k}})=1$ and that $G^{\vee}$ is simply connected. The geometric space of parameters consists of irreducible $G^{\vee}(s)$-equivariant perverse sheaves on the complex vector space $\mathfrak{g}^{\vee}_{q}$. The connected reductive group $G^{\vee}(s)$ acts with finitely many orbits on $\mathfrak{g}^{\vee}_{q}$. This implies that $G^{\vee}(s)$ also acts with finitely many orbits on the dual vector space $(\mathfrak{g}^{\vee}_{q})^{*}$. Identify $(\mathfrak{g}^{\vee}_{q})^{*}$ with $\mathfrak{g}^{\vee}_{q^{-1}}$ by using a fixed nondegenerate invariant symmetric form on $\mathfrak{g}^{\vee}$. Pyasetsky’s duality [Ze] says that there exists a one-to-one correspondence between the orbits in $\mathfrak{g}^{\vee}_{q}$ and in $\mathfrak{g}^{\vee}_{q^{-1}}$, ${{\mathcal{O}}}^{\vee}\to P({{\mathcal{O}}}^{\vee})$, given as follows. For each ${{\mathcal{O}}}^{\vee}\subset\mathfrak{g}^{\vee}_{q}$, $P({{\mathcal{O}}}^{\vee})\subset\mathfrak{g}^{\vee}_{q^{-1}}$ is the unique orbit such that the conormal bundles have the same closures
$\overline{T_{{{\mathcal{O}}}^{\vee}}\mathfrak{g}^{\vee}_{q}}=\overline{T_{P({{\mathcal{O}}}^{\vee})}\mathfrak{g}^{\vee}_{q^{-1}}}$
in $\mathfrak{g}^{\vee}_{q}\times\mathfrak{g}^{\vee}_{q^{-1}}$.
Let $\operatorname{FT}$ denote the Fourier transform from $G^{\vee}(s)$-equivariant perverse sheaves on $\mathfrak{g}^{\vee}_{q}$ to $G^{\vee}(s)$-equivariant perverse sheaves on $\mathfrak{g}^{\vee}_{q^{-1}}$, [EM, §3.2]. Let $\mathsf{IC}({{\mathcal{O}}},{{\mathcal{L}}})$ denote the simple perverse sheaf given by the intersection cohomology complex defined by the equivariant irreducible local system ${{\mathcal{L}}}$ supported on ${{\mathcal{O}}}$. Then $\operatorname{FT}(\mathsf{IC}({{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee}))=\mathsf{IC}({{\mathcal{O}}}^{\vee}_{\mathsf{FT}},{{\mathcal{L}}}^{\vee}_{\mathsf{FT}})$ for an orbit ${{\mathcal{O}}}^{\vee}_{\mathsf{FT}}$ in $\mathfrak{g}^{\vee}_{q^{-1}}$ and an equivariant local system ${{\mathcal{L}}}^{\vee}_{\mathsf{FT}}$. If $\mathfrak{g}^{\vee}=\mathfrak{gl}(n)$, so that all local systems that appear are trivial, [EM, Proposition 7.2] shows that
${{\mathcal{O}}}^{\vee}_{\mathsf{FT}}=P({{\mathcal{O}}}^{\vee}).$
In general however, this is not the case, as one can see already for $G^{\vee}=Sp(4)$ ([EM, Remark 7.5]). We identify, as we may, the $G^{\vee}(s)$-orbits on $\mathfrak{g}^{\vee}_{q^{-1}}$ with the $G^{\vee}(s^{-1})$-orbits on $\mathfrak{g}^{\vee}_{q}$. Then the main result (Theorem 0.1) in loc. cit. is that
${\mathsf{AZ}}(\pi(s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee}))^{*}=\pi(s^{-1},{{\mathcal{O}}}^{\vee}_{\mathsf{FT}},{{\mathcal{L}}}^{\vee}_{\mathsf{FT}}),$ (1.2)
where * in the left hand side denotes the contragredient. In our conjectures below, we will be interested in the nilpotent part of the Langlands parameter of ${\mathsf{AZ}}(\pi)$, that is in $\mathbb{O}^{\vee}_{\operatorname{FT}}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{FT}}.$
1.3. Spaltenstein and Lusztig-Sommers dualities
Let $G$ denote a complex group with maximal torus $T$ with dual $G^{\vee}$ and dual maximal torus $T^{\vee}$. Let $W$ denote the Weyl group of $(G,T)$ and $W^{\vee}$ the Weyl group of $(G^{\vee},T^{\vee})$. Of course, $W\cong W^{\vee}$. Let ${{\mathcal{U}}}_{G}$ and ${{\mathcal{U}}}_{G^{\vee}}$ denote the sets of unipotent elements in $G$ and $G^{\vee}$ respectively. Let $G\backslash{{\mathcal{U}}}_{G}$, $G^{\vee}\backslash{{\mathcal{U}}}_{G^{\vee}}$ denote the corresponding sets of conjugacy classes.
Let $d:G\backslash{{\mathcal{U}}}_{G}\to G^{\vee}\backslash{{\mathcal{U}}}_{G^{\vee}}$ and $d^{\vee}:G^{\vee}\backslash{{\mathcal{U}}}_{G^{\vee}}\to G\backslash{{\mathcal{U}}}_{G}$ denote the duality maps defined by Spaltenstein [Sp]. The images of $d,d^{\vee}$ consist of the set of special unipotent conjugacy classes.
Let $\mathfrak{g}$ be the Lie algebra of $G$. We will also make use of the duality map $d_{S}:G\backslash G\to G^{\vee}\backslash{{\mathcal{U}}}_{G^{\vee}}$ from conjugacy classes of $G$ to unipotent conjugacy classes of $G^{\vee}$, defined by Lusztig [Lu1, §13.3] in the case of special conjugacy classes and generalized by Sommers [So]. Suppose $x\in G$ has a Jordan decomposition $x=su$ and assume $s\in T$. Let $W(s)$ denote the centralizer of $s$ in $W$, this is the Weyl group of the centralizer $G(s)$. Let $E^{W(s)}_{u,1}$ denote the Springer $W(s)$-representation attached to the trivial local system for the unipotent class of $u$ in $G(s)$. Consider the induced representation $\operatorname{Ind}_{W(s)}^{W}(E^{W(s)}_{u,1})$. There exists a unique irreducible $W$-representation $\hat{E}$ in this induced representation such that $b_{\hat{E}}=b_{E^{W(s)}_{u,1}}$, where $b$ denotes the lowest harmonic degree. Regard $\hat{E}$ as an irreducible representation of $W^{\vee}$. By Springer’s correspondence for $G^{\vee}$, there exists a unique unipotent conjugacy class $u^{\vee}$ in $G^{\vee}$ such that $\hat{E}$ is realized in the top cohomology of the Springer fiber for $u^{\vee}$. Define $d_{S}(x)=u^{\vee}$. We recall that
$d_{S}|_{G\backslash{{\mathcal{U}}}_{G}}=d.$
We will often identify the unipotent classes in $G^{\vee}$ with the nilpotent orbits in $\mathfrak{g}^{\vee}$ and think of the image of the map $d_{S}$ as lying in the nilpotent cone of $\mathfrak{g}^{\vee}$.
One can flip the roles of $G$ and $G^{\vee}$ and have the dual map $d^{\vee}_{S}:G^{\vee}\to G$.
1.4. Unramified wavefront set
If $\pi$ is an admissible ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representation, let ${}^{K}\!\mathsf{WF}(\pi)$ denote the unramified wavefront set of $\pi$ as in [Ok]. This is a (finite) collection of nilpotent ${\mathbf{G}}^{\omega}(K)$-orbits in $\mathbf{\mathfrak{g}}^{\omega}(K)$. Using the parametrization in loc. cit., there is a one-to-one correspondence between the set of nilpotent ${\mathbf{G}}^{\omega}(K)$-orbits and pairs
$({{\mathcal{O}}},C):\ {{\mathcal{O}}}\text{ nilpotent ${\mathbf{G}}(\bar{\mathsf{k}})$-orbit in }\mathfrak{g}(\bar{\mathsf{k}}),\ C\text{ conjugacy class in }A({{\mathcal{O}}}),$
where $A({{\mathcal{O}}})$ is a group of components of the centralizer $C_{{\mathbf{G}}(\bar{\mathsf{k}}}(u)$ in ${\mathbf{G}}(\bar{\mathsf{k}})$ of a representative $u$ of ${{\mathcal{O}}}$. By abuse of notation, we think of ${}^{K}\!\mathsf{WF}(\pi)$ as a collection of such pairs $({{\mathcal{O}}},C)$. For each ${{\mathcal{O}}}$ choose a representative $u\in{{\mathcal{O}}}$. For each $C$, we may choose a semisimple representative $\tau$ in $C_{{\mathbf{G}}(\bar{\mathsf{k}})}(u)$. Choosing an isomorphism $\bar{k}\cong\mathbb{C}$, we may identify ${\mathbf{G}}(\bar{\mathsf{k}})$ with the complex group $G={\mathbf{G}}(\mathbb{C})$ with the same root datum, and think of the duality maps in section 1.3 as maps between ${\mathbf{G}}(\bar{\mathsf{k}})$ and $G^{\vee}$. Write
$d_{S}({{\mathcal{O}}},C)=d_{S}(\tau u),$
which is independent of the choice of $u$ and $\tau$. Also write
$d_{S}({}^{K}\!\mathsf{WF}(\pi))=\bigcup_{({{\mathcal{O}}},C)\in{}^{K}\!\mathsf{WF}(\pi)}d_{S}({{\mathcal{O}}},C),$
a union of nilpotent orbits in $\mathfrak{g}^{\vee}$. Define $d({}^{K}\!\mathsf{WF}(\pi))$ similarly.
Also in [Ok], the notion of canonical unramified wavefront set ${}^{K}\!\underline{\mathsf{WF}}(\pi)$ was introduced. This is a collection of classes under a certain equivalence in ${}^{K}\!\mathsf{WF}(\pi)$. The invariant ${}^{K}\!\underline{\mathsf{WF}}(\pi)$ is a set of pairs
$({{\mathcal{O}}},\bar{C}):\ {{\mathcal{O}}}\text{ nilpotent $G$-orbit in }\mathfrak{g},\ \bar{C}\text{ conjugacy class in }\bar{A}({{\mathcal{O}}}),$
where $\bar{A}({{\mathcal{O}}})$ is Lusztig’s canonical quotient. The duality $d_{S}$ factors through the canonical quotient, so we may talk about $d_{S}({}^{K}\!\underline{\mathsf{WF}}(\pi))$.
1.5. Conjectures
Let ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ denote the geometric wavefront set of an admissible representation of ${\mathbf{G}}^{\omega}({\mathsf{k}})$. This is a collection of adjoint nilpotent orbits in $\mathfrak{g}(\bar{\mathsf{k}})$. The nilpotent orbits in $\mathfrak{g}(\bar{\mathsf{k}})$ are ordered with respect to the partial order given by the Zariski closure. The nilpotent orbits in ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ are by definition non-comparable in the closure order. In almost all known examples, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ consists of a single nilpotent orbit if $\pi$ is irreducible, but not always as recent examples of Tsai [Ts] have shown.
If $\mathcal{S}_{1},\mathcal{S}_{2}$ are two sets of nilpotent orbits, we write
$\mathcal{S}_{1}\leq\mathcal{S}_{2},$
if for every $\mathcal{O}_{1}\in\mathcal{S}_{1}$, there exists $\mathcal{O}_{2}\in\mathcal{S}_{2}$ such that $\mathcal{O}_{1}\leq\mathcal{O}_{2}$ in the closure order.
Conjecture 1.1.
Let $\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ be an irreducible ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representation in the L-packet $\Pi(\varphi)=\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})$. Then
  1. (1)
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))\leq d^{\vee}(G^{\vee}\cdot{{\mathcal{O}}}^{\vee}).$ Equivalently,
  2. (1’)
    ${{\mathcal{O}}}^{\vee}\subset\overline{d({}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi)))}$.
Of course, the conjecture is contingent on the existence of a local Langlands correspondence.
Remark 1.2.
While this paper was being finalized, we learned that Conjecture 1.1, as well as the connection with Conjecture 1.9 and Jiang’s conjecture, have also been formulated independently and studied for classical groups by Hazeltine, Liu, Lo, and Shahidi [HLLS].
In terms of the discussion around (1.2), the conjecture could be rephrased as
${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq d^{\vee}(\mathbb{O}^{\vee}_{\mathsf{FT}}),$ (1.3)
whenever we know (or expect) that the action of ${\mathsf{AZ}}$ on the space of geometric parameter is given by the geometric Fourier transform $\operatorname{FT}$. This is the case for unipotent representations.
Remark 1.3.
The conjecture aims to generalize the recent results for Lusztig unipotent representations (trivial $\varphi(I_{\mathsf{k}})$) with real infinitesimal character, see [CMBO1], and for unipotent supercuspidal representations [CMBO2]. In these cases, the upper bound is an equality.
But even in the case of unipotent representations there is an example for the split group $E_{7}$ [CMBO1, §1] that shows that for unipotent representations with nonreal infinitesimal character, the upper bound can be strict.
At the end of section 5, we will comment on other possible versions of the conjecture involving the duality $d_{S}$ and ${}^{K}\!\underline{\mathsf{WF}}(\pi)$.
For ${\mathsf{GL}}(n)$, Conjecture 1.1 is true, in fact a stronger statement holds.
Theorem 1.4 (Mœglin-Waldspurger [MW, II.2]).
For every irreducible ${\mathsf{GL}}(n,{\mathsf{k}})$-representation $\pi$,
${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=d^{\vee}(\mathbb{O}_{\pi}^{\vee}),$ (1.4)
where ${\mathbb{O}}_{\pi}^{\vee}$ is the $G^{\vee}$-saturation of the nilpotent Langlands parameter of $\pi$.
To exemplify Conjecture 1.1, we will verify it for the depth-zero supercuspidal representations of classical groups, using the Langlands parametrization of Lust-Stevens [LuSt], and of $G_{2}$, using Aubert-Xu [AX].
Remark 1.5.
If the representation $\pi(s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ is generic, then its geometric wavefront set is the principal orbit. For the conjecture to hold, we would need $\mathbb{O}^{\vee}_{\mathsf{FT}}=0$, so ${{\mathcal{O}}}^{\vee}$ would need to be the open dense orbit in ${\mathfrak{h}}^{\vee}_{q}$. Moreover, when $H^{\vee}(s)$ is connected, ${{\mathcal{L}}}^{\vee}$ would have to be the trivial local system.
In the case of supercuspidal representations, Conjecture 1.1 takes a simpler, more concrete form. If $\pi$ is supercuspidal, the geometric Langlands parameter is expected to satisfy
  1. (1)
    ${{\mathcal{O}}}^{\vee}={{\mathcal{O}}}_{s}^{\vee}$, the unique open dense $H^{\vee}(s)$-orbit in ${\mathfrak{h}}^{\vee}_{q}$ (this is the case whenever the representation $\pi$ is tempered).
  2. (2)
    In addition, it is immediate from the definition that
    ${\mathsf{AZ}}(\pi)=\pi.$ (1.5)
Conjecture 1.6 (Supercuspidals).
Let $\pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{s},{{\mathcal{L}}}^{\vee})$ be an irreducible supercuspidal ${\mathbf{G}}^{\omega}({\mathsf{k}})$-representation in the L-packet $\Pi(\varphi)=\Pi(s,{{\mathcal{O}}}^{\vee}_{s})$. Set ${\mathbb{O}}^{\vee}_{s}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{s}$. Then
  1. (1)
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq d^{\vee}({\mathbb{O}}^{\vee}_{s}).$
    Equivalently, ${\mathbb{O}}^{\vee}_{s}\leq d({}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi))$.
  2. (2)
    If in addition, ${\mathbf{G}}^{\omega}({\mathsf{k}})$ is split and $\pi$ is depth-zero compactly induced from a cuspidal representation of the maximal hyperspecial parahoric subgroup, then
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=d^{\vee}({\mathbb{O}}^{\vee}_{s}).$
The first part of the conjecture is clearly just a particular case of Conjecture 1.1.
Remark 1.7.
Notice that Conjecture 1.6 is vacuously true for the nonsingular supercuspidal representations, in the sense of Kaletha [Kal], because ${\mathbb{O}}^{\vee}_{s}=0$ so $d^{\vee}({\mathbb{O}}^{\vee}_{s})$ is the principal geometric orbit. In this case, the upper bound is attained only for the generic constituents of the L-packets. In section 3.1, we will compute the wavefront sets of the regular depth-zero supercuspidals parameterized by Kazhdan-Varshavsky [KV] and DeBacker-Reeder [DBR].
In section 4, we will prove
Theorem 1.8 (Depth-zero supercuspidals of classical groups).
Suppose ${\mathbf{G}}({\mathsf{k}})$ is a symplectic, special orthogonal, or unitary group. Conjecture 1.6 holds for all depth-zero supercuspidal representations, using the Langlands parametrization of Lust-Stevens [LuSt].
We mention that Conjectures 1.1 and 1.6 are compatible with a conjecture about Arthur packets formulated by D. Jiang. Suppose $\varphi$ is a tempered L-parameter. Then ${\mathsf{AZ}}(\Pi(\varphi))$ is expected to be an Arthur packet. Let $\pi^{\prime}\in{\mathsf{AZ}}(\Pi(\varphi))$ be one of these “anti-tempered” representations. Then $\pi^{\prime}={\mathsf{AZ}}(\pi(\lambda,s,{{\mathcal{O}}}^{\vee}_{s},{{\mathcal{L}}}^{\vee}))$ for some local system ${{\mathcal{L}}}^{\vee}$. Conjecture 1.1 then says that we should have
${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi^{\prime})\leq d^{\vee}(G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{s}).$ (1.6)
The orbit $G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{s}$ is precisely the “Arthur $SL(2)$${\mathbb{O}}^{\vee}_{A,\varphi}$ attached to the Arthur packet ${\mathsf{AZ}}(\Pi(\varphi))$. In other words, the expectation is
${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi^{\prime})\leq d^{\vee}({\mathbb{O}}^{\vee}_{A,\varphi}),$ (1.7)
for all $\pi^{\prime}\in{\mathsf{AZ}}(\Pi(\varphi))$. This is known to hold (with equality) if $\varphi$ is a “real” unramified L-parameter, see [CMBO3]. Jiang’s conjecture [Ji, Conjecture 4.2] says that this inequality holds for all Arthur packets, when they are defined (not just the above anti-tempered packets) and moreover that there exists one constituent of the Arthur packet where the equality is achieved. One can naturally reformulate Conjecture 1.1 in this case as follows:
Conjecture 1.9.
Let $\pi(\lambda,s,{{\mathcal{O}}}^{\vee},{{\mathcal{L}}}^{\vee})$ be an irreducible representation in the tempered L-packet $\Pi(\varphi)=\Pi(\lambda,s,{{\mathcal{O}}}^{\vee})$. Then
  1. (1)
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))\leq d^{\vee}(G^{\vee}\cdot{{\mathcal{O}}}^{\vee}).$
  2. (2)
    Equality in (1) is achieved for some $\pi$ in the L-packet.
In section 5, we will verify Conjecture 1.9 for $G_{2}$, using the Langlands correspondences from [AX, GS1, GS2] and the wavefront set results of [JLS, LoSa].
In section 2, we will explain that Conjecture 1.9 (via the usual desiderata of Langlands correspondence) implies Conjecture 1.1.
Remark 1.10.
We look at the two extreme cases of Conjectures 1.1 and 1.9.
  1. (a)
    Firstly, let us call an irreducible representation $\pi$ anti-generic if ${\mathsf{AZ}}(\pi)$ is generic. If $\pi$ is anti-generic, then ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=\mathcal{O}_{\textup{reg}}$, the regular nilpotent orbit. The only way that the upper bound conjecture could hold then is if the Langlands nilpotent parameter of $\pi$ is 00. In other words, for every anti-generic representation the image of $\mathbb{C}\subset W_{\mathsf{k}}^{\prime}$ under the Langlands parameter is trivial.
    If $\pi$ is unipotent anti-generic, then it is known that it is spherical (with respect to the maximal special parahoric) and the desired condition on the Langlands parameter holds.
    If $\pi$ is supercuspidal, since ${\mathsf{AZ}}(\pi)=\pi$, the conjecture implies that every generic supercuspidal representation has trivial nilpotent Langlands parameter.
  2. (b)
    Secondly, suppose $\mathcal{O}^{\vee}=0$, so $d^{\vee}(G^{\vee}\cdot\mathcal{O}^{\vee})=\mathcal{O}_{\textup{reg}}.$ Then Conjecture 1.9(1) is trivially satisfied, but (2) says that there should exist a representation $\pi$ in the tempered L-packet such that ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=\mathcal{O}_{\textup{reg}}$. In other words, every non-singular tempered L-packet should contain an anti-generic representation.
Acknowledgements. This paper builds on the previous work of the first-named author with Lucas Mason-Brown and Emile Okada. It is a pleasure to thank them for many useful conversations about the wavefront set. The second-named author would like to thank George Lusztig and Cheng-Chiang Tsai for insightful discussions. We also thank Baiying Liu and Freydoon Shahidi for sharing their work with us and for an interesting correspondence. D.C. was partially supported by the EPSRC grant “New Horizons” EP/V046713/1, 1/2021-12/2023.
2. Parabolic induction
In this section, we discuss the compatibility of our conjectures with parabolic induction. Let $\mathbf{P}=\mathbf{M}\mathbf{N}$ be a $\mathsf{k}$-rational parabolic subgroup of ${\mathbf{G}}$ with a $\mathsf{k}$-rational Levi subgroup $M=\mathbf{M}({\mathsf{k}})$.
If $\pi=\operatorname{Ind}_{\mathbf{P}({\mathsf{k}})}^{{\mathbf{G}}({\mathsf{k}})}(\pi_{M})$ is a parabolically induced representation, where $\pi_{M}$ is an admissible representation of $M$, then the relation between the wavefront sets is well known, see Waldspurger [Wa]:
${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\mathsf{ind}_{\mathbf{M}}^{{\mathbf{G}}}({}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi_{M})),$ (2.1)
where $\mathsf{ind}_{\mathbf{M}}^{{\mathbf{G}}}$ denotes the Lusztig-Spaltenstein induction of nilpotent orbits.
We need to use several known facts about the induction of nilpotent orbits.
Lemma 2.1.
  1. (1)
    If $\mathcal{O}_{1}\leq\mathcal{O}_{2}$ are two nilpotent $\mathbf{M}$-orbits in $\mathfrak{m}$, then $\mathsf{ind}_{\mathbf{M}}^{\mathbf{G}}(\mathcal{O}_{1})\leq\mathsf{ind}_{\mathbf{M}}^{\mathbf{G}}(\mathcal{O}_{2}).$
  2. (2)
    If $\mathcal{O}_{1}^{\vee}\leq\mathcal{O}_{2}^{\vee}$ are two nilpotent $G^{\vee}$-orbits in $\mathfrak{g}^{\vee}$, then $d^{\vee}(\mathcal{O}_{2}^{\vee})\leq d^{\vee}(\mathcal{O}_{1}^{\vee})$.
  3. (3)
    If $\mathcal{O}^{\vee}$ is a nilpotent $M^{\vee}$-orbits in $\mathfrak{m}^{\vee}$, then $d^{\vee}(G^{\vee}\cdot\mathcal{O}^{\vee})=\mathsf{ind}_{\mathbf{M}}^{\mathbf{G}}(d^{\vee}(\mathcal{O}^{\vee})).$
Proof.
Claim (1) is immediate from the definition, or for example from the formula $\overline{\mathsf{ind}_{\mathbf{M}}^{\mathbf{G}}(\mathcal{O})}=\mathbf{G}\cdot(\overline{\mathcal{O}}+\mathfrak{n})$, see [CM, proof of Theorem 7.1.3].
Claims (2) and (3) are in [BV, Proposition A2 (a),(c)]. ∎
Proposition 2.2.
Conjecture 1.9(1) implies Conjecture 1.1.
Proof.
We need to prove that if
${}^{\bar{\mathsf{k}}}\!\mathsf{WF}({\mathsf{AZ}}(\pi))\leq d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$
holds for all irreducible tempered representations of reductive groups, then it holds for all irreducible representations. Here ${\mathbb{O}}^{\vee}_{\pi}$ denotes the nilpotent $G^{\vee}$-orbit of the Langlands parameter of $\pi$. By the parabolic part of the Langlands classification, every irreducible $\pi$ is a quotient of a standard module $i(\mathbf{P},\sigma_{\nu})=i_{\mathbf{P}({\mathsf{k}})}^{\mathbf{G}({\mathsf{k}})}(\sigma_{\nu})$, $\sigma_{\nu}=\sigma\otimes\nu$, where $\sigma$ is an irreducible tempered $\mathbf{M}({\mathsf{k}})$-representation and $\nu$ is an unramified character of $\mathbf{M}({\mathsf{k}})$.
Since the involution ${\mathsf{AZ}}$ commutes with parabolic induction [Au, Theorem 1.7], i.e., in the Grothendieck group
${\mathsf{AZ}}(i_{\mathbf{P}({\mathsf{k}})}^{\mathbf{G}({\mathsf{k}})}(\sigma_{\nu}))=i_{\mathbf{P}({\mathsf{k}})}^{\mathbf{G}({\mathsf{k}})}({\mathsf{AZ}}(\sigma)_{\nu}),$
it follows that ${\mathsf{AZ}}(\pi)$ is a subquotient of $i_{\mathbf{P}({\mathsf{k}})}^{\mathbf{G}({\mathsf{k}})}({\mathsf{AZ}}(\sigma)_{\nu})$. By the definition of the wavefront set then
${}^{\bar{\mathsf{k}}}\!\mathsf{WF}({\mathsf{AZ}}(\pi))\leq{}^{\bar{\mathsf{k}}}\!\mathsf{WF}(i_{\mathbf{P}({\mathsf{k}})}^{\mathbf{G}({\mathsf{k}})}({\mathsf{AZ}}(\sigma)_{\nu})={}^{\bar{\mathsf{k}}}\!\mathsf{WF}(i_{\mathbf{P}({\mathsf{k}})}^{\mathbf{G}({\mathsf{k}})}({\mathsf{AZ}}(\sigma))),$
where the equality comes from the fact that the local character expansion is valid on neighborhoods of compact elements, and so it does not see twists by unramified characters. By (2.1), it follows that
${}^{\bar{\mathsf{k}}}\!\mathsf{WF}({\mathsf{AZ}}(\pi))\leq\mathsf{ind}_{\mathbf{M}}^{{\mathbf{G}}}({}^{\bar{\mathsf{k}}}\!\mathsf{WF}({\mathsf{AZ}}(\sigma))).$
Now assuming Conjecture 1.9 for tempered modules, we have ${}^{\bar{\mathsf{k}}}\!\mathsf{WF}({\mathsf{AZ}}(\sigma))\leq d^{\vee}_{M^{\vee}}({\mathbb{O}}^{\vee}_{\sigma})$. Then Lemma 2.1(1),(3) implies
${}^{\bar{\mathsf{k}}}\!\mathsf{WF}({\mathsf{AZ}}(\pi))\leq\mathsf{ind}_{\mathbf{M}}^{{\mathbf{G}}}(d^{\vee}_{M^{\vee}}({\mathbb{O}}^{\vee}_{\sigma}))=d^{\vee}(G^{\vee}\cdot{\mathbb{O}}^{\vee}_{\sigma}).$
Recall that $\pi$ is the Langlands quotient of $i(\mathbf{P},\sigma_{\nu})$, hence by the desiderata of the Langlands correspondence, ${\mathbb{O}}^{\vee}_{\pi}=G^{\vee}\cdot{\mathbb{O}}^{\vee}_{\sigma}.$
3. Depth-zero supercuspidal representations
[…]
The finite analogue of the wavefront set for $\sigma$ is the Kawanaka wavefront set $\mathsf{WF}(\sigma)$, [Kaw].
Theorem 3.1 ([CMBO2, Theorem 3.0.4]).
The canonical unramified wavefront set of $\pi(x,\sigma)$ is
${}^{K}\!\underline{\mathsf{WF}}(\pi(x,\sigma))=(\mathsf{WF}(\sigma),x).$
Lusztig gave an explicit description of $\mathsf{WF}(\mu)$, $\mu$ an irreducible representation of ${\mathbf{G}}^{\omega}_{x}({\mathbb{F}}_{q})$ in terms of his parametrization [Lu1, Theorem 4.23]. Let $({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})$ be the dual finite reductive group. Then $\mu$ is parametrized by a triple $(\tau,u_{0}^{\vee},y)$, where $\tau$ is a semisimple element in $({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})$, $u_{0}^{\vee}$ is a special unipotent element in the centralizer $C_{({\mathbf{G}}^{\omega}_{x})^{\vee}({\mathbb{F}}_{q})}(\tau)$ and $y$ is an element of the unipotent family $C_{u_{0}^{\vee}}$ parametrized by $u_{0}^{\vee}$. The element $y$ will not be important for us. Let $\widetilde{C}_{u_{0}^{\vee}}=C_{\widetilde{u}_{0}^{\vee}}$ be the family obtained by applying Alvis-Curtis duality to $C_{u_{0}^{\vee}}$. Let $d_{S}^{\vee}$ be the duality map in this finite field setting, this is defined exactly as above, see [Lu1, §13.3].
Theorem 3.2 ([Lu2, Theorem 11.2]).
With the above notation, $\mathsf{WF}(\mu)=d_{S}^{\vee}(\tau\widetilde{u}_{0}^{\vee})$. If $\mu=\sigma$ is cuspidal, then $\mathsf{WF}(\sigma)=d_{S}^{\vee}(\tau u_{0}^{\vee})$.
This means that the canonical unramified wavefront set of a supercuspidal representation $\pi(x,\sigma)$ is
${}^{K}\!\underline{\mathsf{WF}}(\pi(x,\sigma))=(d_{S}^{\vee}(\tau u_{0}^{\vee}),x).$ (3.1)
Since we are in a depth-zero case, the Langlands parameter $\varphi$ is trivial on the wild inertia group, and the image $\varphi(I_{\mathsf{k}})$ is a cyclic group generated by an element $\lambda\in{}^{L}\!G$ of finite order prime to $q$. Hence ${\mathfrak{h}}^{\vee}={\mathfrak{g}^{\vee}}^{\lambda}$ is a pseudo-Levi subalgebra of $\mathfrak{g}^{\vee}$. Conjecture 1.6 is equivalent to
$G\cdot d_{S}^{\vee}(\tau u_{0}^{\vee})\leq d^{\vee}(G^{\vee}\cdot{{\mathcal{O}}}_{s}^{\vee}).$ (3.2)
We do not know a uniform way to attach the Langlands parameter to a depth-zero supercuspidal representation, so we will only verify the conjecture in a number of cases that exist in the literature.
3.1. Regular supercuspidal representations
For all regular depth-zero supercuspidal representations [KV, DBR], the conjecture is trivially true since in that case ${\mathfrak{h}}^{\vee}$ is just a Cartan subalgebra of $\mathfrak{g}^{\vee}$, so the only possible Langlands nilpotent element is 00. Hence the conjecture just says that ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq{{\mathcal{O}}}_{p}$, where ${{\mathcal{O}}}_{p}$ is the principal ${\mathbf{G}}(\bar{\mathsf{k}})$-orbit, which is true, but trivial. On the other hand, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)={{\mathcal{O}}}_{p}$ if and only if $\pi$ is a generic representation, and this is not always the case. In general, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ can be equal to the ${\mathbf{G}}(\bar{\mathsf{k}})$-saturation of the principal nilpotent orbit in a maximal pseudo-Levi, as we explain next.
Let $I_{t}=I_{\mathsf{k}}/I^{+}_{\mathsf{k}}$ be the tame quotient of the inertia group, where $I^{+}_{\mathsf{k}}$ is the wild inertia subgroup. The depth-zero condition means that the parameter $\lambda$ in the introduction factors through $I_{t}$. Write the L-group as ${}^{L}G=G^{\vee}\rtimes\langle\theta^{\vee}\rangle$, where $\theta^{\vee}$ is a finite-order automorphism giving the inner class of ${\mathbf{G}}$. Suppose $G^{\vee}$ is semisimple. In the terminology of DeBacker and Reeder, a tame regular semisimple elliptic Langlands parameter is given a $W^{\vee}$-conjugacy class of pairs $(\lambda,w)$, where:
  1. (1)
    a continuous homomorphism $\lambda:I_{t}\to T^{\vee}$ such that $Z_{G^{\vee}}(\lambda)=T^{\vee}$, and
  2. (2)
    a Weyl group element $w\in W^{\vee}=N_{G^{\vee}}(T^{\vee})/T^{\vee}$ such that $w^{\vee}\circ\theta^{\vee}\circ\lambda^{q}=\lambda$ and ${T^{\vee}}^{w^{\vee}\theta^{\vee}}$ is finite.
The element $w^{\vee}\theta^{\vee}$ is the image of the inverse Frobenius under the Langlands homomorphism $\varphi$. Then a complete (enhanced) Langlands parameter is given by a triple $(\lambda,w,\rho)$, where $\rho$ is an irreducible representation of the finite group ${T^{\vee}}^{w^{\vee}\theta^{\vee}}=Z_{G^{\vee}}(\varphi).$ If $X$ is the weight lattice of ${\mathbf{G}}$, and $W$ is the Weyl group of ${\mathbf{G}}$, $\theta$ the Galois automorphism giving $\theta^{\vee}$, there is a natural isomorphism
$X/(1-w\theta)X\longleftrightarrow\mathsf{Irr}~{}{T^{\vee}}^{w^{\vee}\theta^{\vee}},\quad x\mapsto\rho_{x}.$ (3.3)
By abuse of notation, denote $x$ also the preimage in $X$ of an element in $X/(1-w\theta)X$. Let $\pi(\lambda,w,0,x)$ denote the supercuspidal representation attached to $(\lambda,w,x)$ in loc. cit. The 00 here stands for the zero nilpotent orbit, i.e., the image of the $SL(2)$ under the Langlands parameter must be trivial (since ${\mathfrak{g}^{\vee}}^{\lambda(I_{\mathsf{k}})}=\mathfrak{t}^{\vee}$).
Proposition 3.3.
The geometric wavefront set ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi(\lambda,w,0,x))$ is the ${\mathbf{G}}(\bar{\mathsf{k}})$-saturation of the principal nilpotent orbit of the maximal pseudo-Levi subgroup $Z_{{\mathbf{G}}(\bar{\mathsf{k}})}(x).$
Proof.
The claim follows at once, after we recall the construction of $\pi(\lambda,w,0,x)$. The element $x$ defines a cocycle $u_{x}\in H^{1}({\mathsf{k}},{\mathbf{G}})$ which gives the inner twist $\omega_{x}$ of ${\mathbf{G}}$ and $F_{x}$ a twisted Frobenius map and a maximal parahoric subgroup ${\mathbf{G}}_{x}$. Then ${\mathbf{G}}_{x}^{F_{x}}$ is a maximal parahoric subgroup of ${\mathbf{G}}^{\omega_{x}}({\mathsf{k}})$. The pair $(w,x)$ defines an $F_{x}$-minisotropic torus $T_{x}$ and $\lambda$ defines a depth-zero character $\chi_{x}$ of $T_{x}$. Then Deligne-Lusztig induction defines an irreducible cuspidal representation of the reductive quotient $\overline{{\mathbf{G}}_{x}^{F_{x}}}$
$\kappa_{x}=\pm R_{T_{x}}^{\overline{{\mathbf{G}}_{x}^{F_{x}}}}(\chi_{x}).$
The important thing for us is that this cuspidal representation is generic, in the sense that the Kawanaka wavefront set is principal. The supercuspidal representation is compactly induced
$\pi(\lambda,w,0,x)=\operatorname{ind}_{Z^{F_{x}}{\mathbf{G}}_{x}^{F_{x}}}^{{\mathbf{G}}^{\omega_{x}}({\mathsf{k}})}(\kappa_{x}).$
The wavefront set is given by Theorem 3.1. ∎
Conjecture 1.6 clearly holds. Moreover, so does Conjecture 1.9(2) since the L-packet of supercuspidals contains a generic supercuspidal representation.
4. Depth-zero representations of classical groups
In [LuSt], Lust and Stevens constructed L-packets of depth-zero supercuspidal representations of classical groups. We will verify Conjecture 1.6 for these $L$-packets. Let ${\mathcal{G}}$ be a connected reductive group of classical type with connected center over ${\mathbb{F}}_{q}$, $q$ odd. Let $F$ be the Frobenius map. Let ${\mathcal{G}}^{*}$ be the dual group over ${\mathbb{F}}_{q}$ with dual Frobenius also denoted $F$. Let $\mathsf{Irr}{\mathcal{G}}^{F}$ be the set of equivalence classes of irreducible complex ${\mathcal{G}}^{F}$-representations. According to Lusztig’s classification, we can partition
$\mathsf{Irr}{\mathcal{G}}^{F}=\sqcup_{s}\mathsf{Irr}_{s}{\mathcal{G}}^{F},$
where $s$ runs over the conjugacy classes of semisimple elements in ${{\mathcal{G}}^{*}}^{F}$. The piece $\mathsf{Irr}_{1}{\mathcal{G}}^{F}$ is the set of irreducible unipotent representations. Since $Z({\mathcal{G}})$ is connected, the rational semisimple conjugacy classes in ${\mathcal{G}}^{*}$ are in one-to-one correspondence with the geometric $F$-stable semisimple conjugacy classes in ${\mathcal{G}}^{*}$ [Car, Proposition 3.7.3].
The above decomposition is realized as follows. By [Car, Theorem 4.4.6], there exists a natural one-to-one correspondence between geometric conjugacy classes of pairs $(T,\theta)$, where ${\mathcal{T}}$ is a maximal $F$-stable torus in ${\mathcal{G}}$ and $\theta$ is a character of $T$, and $F$-stable semisimple conjugacy classes in ${\mathcal{G}}^{*}$. Let $s\in{\mathcal{T}}^{*}\subset{\mathcal{G}}^{*}$, then $\mathsf{Irr}_{s}{\mathcal{G}}^{F}$ consists of all the constituents of the induced representations $R_{{\mathcal{T}}}^{{\mathcal{G}}}(s)$ where we think of $s$ as a character of $T$.
Lusztig’s Jordan decomposition says that there is a one-to-one correspondence
$\Psi_{s}^{{\mathcal{G}}}:\mathsf{Irr}_{s}{\mathcal{G}}^{F}\longleftrightarrow\mathsf{Irr}_{1}({\mathcal{G}}_{s}^{F}),$
where ${\mathcal{G}}_{s}$ is the endoscopic group of ${\mathcal{G}}$ with dual ${\mathcal{G}}^{*}_{s}:=Z_{{\mathcal{G}}^{*}}(s)$. To have cuspidal representations, we need $Z({\mathcal{G}}^{*})^{0}$ and $Z({\mathcal{G}}_{s}^{*})^{0}$ to have the same $F$-rank. In this case, $\Psi_{s}^{{\mathcal{G}}}$ gives a correspondence of cuspidal represenattions. In other words, the set of cuspidal ${\mathcal{G}}^{F}$-representations are in one-to-one correspondence with ${\mathcal{G}}^{*}$-pairs $(s,\tau)$ where $s\in{{\mathcal{G}}^{*}}^{F}$ is semisimple of equal rank and $\tau$ is a cuspidal unipotent representation of ${\mathcal{G}}_{s}^{F}$.
For classical groups, the classification of cuspidal unipotent representations shows that there is at most one cuspidal unipotent representation. On the other hand, the conjugacy classes of $s\in{{\mathcal{G}}^{*}}^{F}$ are given by characteristic polynomials $P_{s}\in{\mathbb{F}}_{q}[x]$,
$P_{s}(x)=\prod_{P}P(x)^{a_{P}},$
where the product runs over all irreducible self-dual monic polynomials over ${\mathbb{F}}_{q}$ and nonnegative integers $a_{P}$ such that $\sum_{P}a_{P}\deg(P)=\dim V$, where $V$ is the defining $\mathbb{F}_{q}$-representation of ${\mathcal{G}}^{*}$. In order for the semisimple class to parametrize a cuspidal representation, one needs:
  1. (1)
    if $P(x)\neq x\pm 1$, then $a_{P}=\frac{1}{2}(m_{P}^{2}+m_{P})$ for some integer $m_{P}\geq 0.$
  2. (2)
    if we set $a_{\pm}=a_{x\mp 1}$, there are some conditions on $a_{\pm}$ depending on the type of the group, see [LuSt, §7.2].
Example 4.1.
Suppose ${\mathcal{G}}=Sp(2n)$. If $P(x)$ appears in $P_{s}(x)$ and $\deg(P)>1$, then necessarily $\deg(P)$ is even. We denote $\deg(P)=2k_{P}$, $k_{P}>0$ integer. Otherwise,
$a_{+}=2(m_{+}^{2}+m_{+})+1,\quad a_{-}=2m_{-}^{2},$
for some nonnegative integers $m_{\pm}$. Then a cuspidal representation of $Sp(2n,{\mathbb{F}}_{q})$ is uniquely determined by a collection of nonnegative integers $m_{\pm}$, $m_{P},k_{P}$ such that
$\sum_{\deg(P)>1}k_{P}\frac{m_{P}^{2}+m_{P}}{2}+(m_{+}^{2}+m_{+})+m_{-}^{2}=n.$ (4.1)
The $\deg(P)>1$ terms, if they appear, correspond to endoscopic unitary groups, whereas the $m_{+}$ term corresponds to an endoscopic $Sp$ subgroup and $m_{-}$ to an endoscopic even orthogonal group.
4.1. Nilpotent Langlands parameters
Let $G$ be a classical group over a $p$-adic field ${\mathsf{k}}$. By this we mean, a symplectic, special orthogonal, or a unitary group. Every depth-zero supercuspidal representation is compactly-induced
$\pi=\text{ind}_{K}^{G}(\tau_{1}\boxtimes\tau_{2}),$ (4.2)
where $K$ is the normalizer of a maximal parahoric subgroup with reductive quotient $\bar{K}={{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}$ over ${\mathbb{F}}_{q}$, and $\tau_{1}\boxtimes\tau_{2}$ is a cuspidal representation of ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})$. The explicit cases are as follows [LuSt, §2], for more details about the notation, see loc. cit.:
  1. (1)
    $G$ symplectic, ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})=Sp(2n_{1},{\mathbb{F}}_{q})\times Sp(2n_{2},{\mathbb{F}}_{q})$;
  2. (2)
    $G$ special orthogonal, ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})=SO(n_{1}+n_{1}^{an},n_{1},{\mathbb{F}}_{q})\times SO(n_{2}+n_{2}^{an},n_{2},{\mathbb{F}}_{q})$, where $SO(m+m^{an},m,{\mathbb{F}}_{q})$ denotes the special orthogonal group with form of Witt index $m$ and anisotropic part of dimension $m^{an}\leq 2$;
  3. (3)
    $G$ unramified unitary, ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}=U(2n_{1}+n_{1}^{an})\times U(2n_{2}+n_{2}^{an})$, a product of two unitary groups over ${\mathbb{F}}_{q}$;
  4. (4)
    $G$ ramified unitary, then ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}$ is
    $SO(n_{1}+n_{1}^{an},n_{1},{\mathbb{F}}_{q})\times Sp(2n_{2},{\mathbb{F}}_{q})\text{ or }Sp(2n_{1},{\mathbb{F}}_{q})\times SO(n_{2}+n_{2}^{an},n_{2},{\mathbb{F}}_{q}).$
Let $(m_{\pm}^{(i)},m_{P}^{(i)},k_{P}^{(i)})$, $i=1,2$ be the collections of integers that characterizes $\tau_{i}$. The nilpotent Langlands parameter of $\pi$ is a nilpotent orbit ${\mathbb{O}}^{\vee}_{\pi}$ in $\mathfrak{g}^{\vee}$, which is either $gl(n,{\mathbb{C}})$, $sp(2n,{\mathbb{C}})$ or $so(m,{\mathbb{C}})$. Via the Jordan normal form, ${\mathbb{O}}^{\vee}_{\pi}$ is parametrized by a partition $\lambda$ of $n$, $2n$, or $m$, respectively. If $\mathfrak{g}^{\vee}=so(m,{\mathbb{C}})$, then the partition is orthogonal: each even part appears with even multiplicity. If $\mathfrak{g}^{\vee}=sp(2n,{\mathbb{C}})$, then the partition is symplectic: each odd part appears with even multiplicity.
Lust and Stevens [LuSt, §8] attach to $\pi$ the nilpotent orbit ${\mathbb{O}}^{\vee}_{\pi}$ given by the partition obtained by the following procedure. For every integer $\alpha$ denote by $[\alpha]$ the string $\{\alpha,\alpha-2,\alpha-4,\dots,1\}$ if $\alpha>0$ is odd, $\{\alpha,\alpha-2,\alpha-4,\dots,2\}$ if $\alpha>0$ is even, and $[\alpha]=\emptyset$ if $\alpha\leq 0$. Start with an empty partition and
  1. (1)
    if $k_{P}>0$, append $2k_{P}$ copies of $[m_{P}^{(1)}+m_{P}^{(2)}]$ and $2k_{P}$ copies of $[|m_{P}^{(1)}-m_{P}^{(2)}|-1]$.
For unramified unitary groups, this completes the nilpotent Langlands parameter. For the other classical groups, we need to append more strings defined in terms of $m_{\pm}^{(i)}$ and the type of the $p$-adic group.
Symplectic groups.
  1. (2s)
    append the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})+1]$ and $[2|m_{+}^{(1)}-m_{+}^{(2)}|-1]$;
  2. (3s)
    append the strings $[2(m_{-}^{(1)}+m_{-}^{(2)})-1]$ and $[2|m_{-}^{(1)}-m_{-}^{(2)}|-1]$.
Ramified unitary groups. Without loss of generality, assume ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})=Sp(2n_{1},{\mathbb{F}}_{q})\times SO(n_{2}+n_{2}^{an},n_{2},{\mathbb{F}}_{q})$. There are two cases depending on the parity of $n_{2}^{an}$:
  1. (2r)
    append the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})+1]$ and $[2|m_{+}^{(1)}-m_{+}^{(2)}|-1]$ if $n_{2}^{an}=1$, or the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})]$ and $[2|m_{+}^{(1)}-m_{+}^{(2)}+1|-1]$ if $n_{2}^{an}$ is even;
  2. (3r)
    append the strings $[2(m_{-}^{(1)}+m_{-}^{(2)})]$ and $[2|m_{-}^{(1)}-m_{-}^{(2)}-1|-1]$ if $n_{2}^{an}=1$, or the strings $[2(m_{-}^{(1)}+m_{-}^{(2)})-1]$ and $[2|m_{-}^{(1)}-m_{-}^{(2)}|-1]$ if $n_{2}^{an}$ is even.
Special orthogonal groups. There are three cases depending on the parities of $n_{1}^{an}$ and $n_{2}^{an}$.
  1. (2o)
    • If $n_{1}^{an}=1$ and $n_{2}^{an}$ is even, append the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})]$ and $[|2(m_{+}^{(1)}-m_{+}^{(2)})+1|-1]$.
    • If $n_{1}^{an}$ and $n_{2}^{an}$ are both even, append the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})-1]$ and $[|2(m_{+}^{(1)}-m_{+}^{(2)})|-1]$.
    • If $n_{1}^{an}=n_{2}^{an}=1$, append the strings $[2(m_{+}^{(1)}+m_{+}^{(2)})+1]$ and $[|2(m_{+}^{(1)}-m_{+}^{(2)})|-1]$.
  2. (3o)
    Exactly the same, but with $m_{-}$ in place of $m_{+}$.
4.2. Symplectic groups
We verify Conjecture 1.6 for the case of depth-zero supercuspidal representations of $G=Sp(2n,{\mathsf{k}})$, by computing the wavefront set and compare it with the nilpotent Langlands parameter defined above.
Definition 4.2.
If $\alpha,\beta$ are two partitions of $m,n$, respectively, the sum $\alpha+\beta$ is a partition of $m+n$ given as follows. Suppose $\alpha=(\alpha_{1},\alpha_{2},\dots,\alpha_{\ell})$ and $\beta=(\beta_{1},\beta_{2},\dots,\beta_{\ell})$ with $0\leq\alpha_{1}\leq\alpha_{2}\leq\dots\leq\alpha_{\ell}$, $0\leq\beta_{1}\leq\beta_{2}\leq\dots\leq\beta_{\ell}$. (By allowing entries of 00, we may assume there is the same number of parts in $\alpha$ and $\beta$.) Then $\alpha+\beta=(\alpha_{1}+\beta_{1},\alpha_{2}+\beta_{2},\dots,\alpha_{\ell}+\beta_{\ell}).$
This operation on partitions plays a key role in the explicit computation of $j$-induction for representations of the Weyl group of classical types.
At one extreme suppose that only endoscopic unitary factors appear, that is only polynomials $P$ with $\deg P>1$. Hence ${\mathbb{O}}^{\vee}_{\pi}$ is obtained by appending $2k_{P}$ copies of $[m_{P}^{(1)}+m_{P}^{(2)}]$ and $2k_{P}$ copies of $[|m_{P}^{(1)}-m_{P}^{(2)}|-1]$, for each $k_{P}>0$. On each $Sp(2n_{i})$, $i=1,2$, factor, we have
$\sum_{\deg P>1}k_{p}a_{p}^{(i)}=n_{i},\quad a_{P}^{(i)}=\frac{{m_{P}^{(i)2}}+m_{P}^{(i)}}{2}.$
To compute the wavefront set, we need to compute the $j$-induction on each factor. To simplify the notation, let us drop temporarily the indices $i$. Let $W_{n}$ be the Weyl group of $G^{\vee}$. Then, the calculation amounts to determining the irreducible $W_{n}$-representation given by
$j_{\prod_{P}S_{a_{P}}^{k_{P}}}^{W_{n}}\left(\boxtimes_{P}(1,2,\dots,m_{P})^{\boxtimes k_{P}}\right).$
Using the formulae for $j$-induction [Lu3, §3.1,§4.4], this equals
$j_{\prod_{P}S_{k_{P}a_{P}}}^{W_{n}}\left(\boxtimes_{P}(k_{P},2k_{P},\dots,m_{P}k_{P})\right)=j_{\prod_{P}W_{k_{P}a_{P}}}^{W_{n}}(\sigma_{P}),$ (4.3)
where
$\sigma_{P}=\begin{cases}(k_{P},3k_{P},\dots,m_{P}k_{P})\times(2k_{P},4k_{P},\dots,(m_{P}-1)k_{P}),&m_{P}\text{ odd}\\ (2k_{P},4k_{P},\dots,m_{P}k_{P})\times(k_{P},3k_{P},\dots,(m_{P}-1)k_{P}),&m_{P}\text{ even}.\end{cases}$
A direct calculation with the algorithm for the Springer correspondence gives that the type $C$ nilpotent orbit for $\sigma_{P}$ is given by the partition
[…]
This is true in both cases, $m_{P}$ even or odd.
Lemma 4.3.
The irreducible $W_{n}$-representation which is $j$-induced
$j_{\prod_{P}S_{a_{P}}^{k_{P}}}^{W_{n}}\left(\boxtimes_{P}(1,2,\dots,m_{P})^{\boxtimes k_{P}}\right)$
is attached in the Springer correspondence to the symplectic nilpotent orbit
$\sum_{P}(2k_{P},4k_{P},\dots,2m_{P}k_{P}),$
where the sum is in the sense of Definition 4.2.
Proof.
By the calculation leading up to (4.4), we need to show that the nilpotent orbit for $j_{\prod_{P}W_{k_{P}a_{P}}}^{W_{n}}(\boxtimes_{P}\sigma_{P})$ is as claimed in the statement of the lemma. Since $j$-induction works in stages, this comes down to the case of two polynomials $P_{1},P_{2}$, that is to the easy verification that $j_{W_{1}\times W_{2}}^{W}(\sigma_{1}\boxtimes\sigma_{2})$ is the Springer representation for the nilpotent $(2k_{1},4k_{1},\dots,2m_{1}k_{1})+(2k_{2},4k_{2},\dots,2m_{2}k_{2})$. ∎
Proposition 4.4.
Let $\pi$ be a depth-zero supercuspidal $Sp(2n,\mathsf{k})$-representation as above. Suppose there are only endoscopic unitary factors, i.e., only polynomials $P$ with $\deg P>1$ appear. Then
  1. (1)
    ${\mathbb{O}}^{\vee}_{\pi}=\bigsqcup_{P}[m_{P}^{(1)}+m_{P}^{(2)}]^{2k_{P}}\sqcup[|m_{P}^{(1)}-m_{P}^{(2)}|-1]^{2k_{P}}$.
  2. (2)
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\bigsqcup_{i=1,2}\left(\sum_{P}(2k_{P},4k_{P},\dots,2m_{P}^{(i)}k_{p})\right)$
  3. (3)
    Conjecture 1.6 holds in this case.
Proof.
Claim (1) is just a restatement of the Lust-Stevens parametrization. Claim (2) is directly implied by Lemma 4.3. It remains to verify (3).
Suppose first that $n_{1}=n$, $n_{2}=0$, this is the hyperspecial case. Write $m_{P}$ for $m_{P}^{(1)}$. Then $\mathbb{O}^{\vee}_{\pi}$ is simply
${\mathbb{O}}^{\vee}_{\pi}=\bigsqcup_{P}(\underbrace{m_{P},\dots,m_{P}}_{2k_{P}},\underbrace{m_{P}-1,\dots,m_{P}-1}_{2k_{P}},\dots,\underbrace{1,\dots,1}_{2k_{P}}),$
while
${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\sum_{P}(2k_{P},4k_{P},\dots,2m_{P}k_{p})$
Notice that
$d^{\vee}((\underbrace{m_{P},\dots,m_{P}}_{2k_{P}},\underbrace{m_{P}-1,\dots,m_{P}-1}_{2k_{P}},\dots,\underbrace{1,\dots,1}_{2k_{P}}))$ $=(\underbrace{m_{P},\dots,m_{P}}_{2k_{P}},\underbrace{m_{P}-1,\dots,m_{P}-1}_{2k_{P}},\dots,\underbrace{1,\dots,1}_{2k_{P}})^{t}$
$=(2k_{P},4k_{P},\dots,2m_{P}k_{p})$
which proves that $d^{\vee}({\mathbb{O}}_{\pi}^{\vee})={}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ if there is a single $P$. If there is more than one polynomial $P$, the claim follows as well by the compatibility
$(\alpha\sqcup\beta)^{t}=\alpha^{t}+\beta^{t}$ (4.5)
for two (or more) partitions $\alpha,\beta$.
Now suppose both $n_{1},n_{2}>0$. The key observation is that for all $k$ and $m^{(1)}\geq m^{(2)}$,
$d^{\vee}([m^{(1)}+m^{(2)}]^{2k}\sqcup[|m^{(1)}-m^{(2)}|-1]^{2k})$ $=\left((m^{(1)}+m^{(2)},m^{(1)}+m^{(2)}-2,\dots,m^{(1)}-m^{(2)}+2\right.,$
$\left.m^{(1)}-m^{(2)},m^{(1)}-m^{(2)}-1,\dots,2,1)^{2k}\right)^{t}$
$=(2k,4k,\dots,2m^{(1)}k)\sqcup(2k,4k,\dots,2m^{(2)}k),$
which proves (3) (with equality in Conjecture 1.6) in the case of a single polynomial $P$.
Finally, in the general case with more polynomials, we use again the compatibility (4.5) and see that
$d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ $=\left(\bigsqcup_{P}[m_{P}^{(1)}+m_{P}^{(2)}]^{2k_{P}}\sqcup[|m_{P}^{(1)}-m_{P}^{(2)}|-1]^{2k_{P}}\right)^{t}$
$=\sum_{P}\left((2k_{P},4k_{P},\dots,2m_{P}^{(1)}k_{P})\sqcup(2k_{P},4k_{P},\dots,2m_{P}^{(2)}k_{P})\right)$
It is easy to see that this is greater in the partition order than ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\bigsqcup_{i=1,2}\left(\sum_{P}(2k_{P},4k_{P},\dots,2m_{P}^{(i)}k_{p})\right)$. Notice that the difference between the two formulae is simply that $\sigma$ and $\sqcup_{i=1,2}$ are interchanged.
To see that the inequality ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)\leq d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ can be strict, consider the case when there are just two distinct polynomials $P_{1},P_{2}$, where $P_{i}$ occurs on the $Sp(2n_{i})$ factor. Then
${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=(2k_{P_{1}},4k_{P_{1}},\dots,2m_{P_{1}}^{(1)}k_{P_{1}})\sqcup(2k_{P_{2}},4k_{P_{2}},\dots,2m_{P_{2}}^{(2)}k_{P_{2}})$
while
$d^{\vee}({\mathbb{O}}^{\vee}_{\pi})=(2k_{P_{1}},4k_{P_{1}},\dots,2m_{P_{1}}^{(1)}k_{P_{1}})+(2k_{P_{2}},4k_{P_{2}},\dots,2m_{P_{2}}^{(2)}k_{P_{2}}).$
The other extreme is that only endoscopic factors of type $B$ and $D$ occur in the construction of the cuspidal representation on the finite reductive group of the parahoric $Sp(2n_{1})\times Sp(2n_{2})$.
Proposition 4.5.
Let $\pi$ be a depth-zero supercuspidal $Sp(2n,\mathsf{k})$-representation as above. Suppose there are no endoscopic unitary factors, i.e., no polynomials $P$ with $\deg P>1$ appear. Then
  1. (1)
    ${\mathbb{O}}^{\vee}_{\pi}=[2(m_{+}^{(1)}+m_{+}^{(2)})+1]\sqcup[2|m_{+}^{(1)}-m_{+}^{(2)}|-1]\sqcup[2(m_{-}^{(1)}+m_{-}^{(2)})-1]\sqcup[2|m_{-}^{(1)}-m_{-}^{(2)}|-1].$
  2. (2)
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\bigsqcup_{i=1,2}[2(m_{+}^{(i)}+m_{-}^{(i)})]\sqcup\begin{cases}[2(m_{+}^{(i)}-m_{-}^{(i)})],&m_{+}^{(i)}\geq m_{-}^{(i)},\\ [2(m_{-}^{(i)}-m_{+}^{(i)}-1)],&m_{+}^{(i)}<m_{-}^{(i)}\end{cases}.$
  3. (3)
    Conjecture 1.6 holds in this case.
Proof.
Claim (1) is just a restatement of the above procedure by Lust-Stevens, listed here for convenience. Claim (2) follows from formula (3.1) by a direct calculation using $j$-induction. More precisely, ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$ is the saturation (concatenation in terms of the partitions) of the two Kawanaka wavefront set nilpotent orbits on the two factors $Sp(2n_{i})$ of the maximal parahoric.
Fix one factor and drop the index $(i)$ from the notation for simplicity. The Kawanaka wavefront set is the symplectic nilpotent orbit for which the Springer representation (attached to the trivial local system) is given the $j$-induced
$j_{W(B)\times{W(D)}}^{W_{n}}\left(((1,2,\dots,m_{+})\times(1,2,\dots,m_{+}))\boxtimes(1,2,\dots,m_{-})\times(1,2,\dots,m_{-}-1))\right),$
where the notation indicates that the first factor is of type $B$ and the second of type $D$. We use the customary bipartition notation $\alpha\times\beta$ for an irreducible representation of the Weyl groups of types $B$ or $D$.
Using the formulae for $j$-induction from [Lu3, §5], we see this $j$-induced representation is the bipartition
$\left((1,2,\dots,m_{+})+(1,2,\dots,m_{-})\right)\times\left((1,2,\dots,m_{+})+(1,2,\dots,m_{-}-1)\right),$
where $+$ has the meaning from Definition 4.2. Applying the combinatorial algorithm for the Springer representation for type $C$, we find that the corresponding symplectic nilpotent orbit is
$[2(m_{+}+m_{-})]\sqcup\begin{cases}[2(m_{+}-m_{-})],&m_{+}\geq m_{-},\\ [2(m_{-}-m_{+}-1)],&m_{+}<m_{-}\end{cases},$
and Claim (2) is proven.
Claim (3) follows by computing Spaltenstein duality $d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ similarly to the proof of the previous proposition. In the case when $n_{2}=0$, so $m_{+}^{(2)}=m_{-}^{(2)}=0$, we get equality $d^{\vee}({\mathbb{O}}^{\vee}_{\pi})={}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)$. ∎
We can now consider the general case by putting together the two cases in Propositions 4.5 and 4.4.
Theorem 4.6.
Conjecture 1.6 holds in the case when $\pi$ be a depth-zero supercuspidal $Sp(2n,\mathsf{k})$-representation.
Proof.
Let $\lambda_{\pm}^{\vee}$ be the orthogonal partition for the dual nilpotent orbit given by the Lust-Stevens algorithm for the polynomials $P$ with degree equal to $1$. That is, $\lambda_{\pm}^{\vee}$ has an expression like in Proposition 4.5(1). Let $\mu_{\pm}$ be the symplectic partition for the wavefront set again for the $\deg P>1$ part as in Proposition 4.5(2).
Similarly, let $\lambda^{\vee}_{\text{uni}}$ and $\mu_{\text{uni}}$ be the orthogonal and symplectic partitions that encode the dual nilpotent orbit and the wavefront set, respectively, in the case when all polynomials have degree strictly greater than $1$, i.e., in the setting of Proposition 4.4(1),(2).
Then
${\mathbb{O}}^{\vee}_{\pi}=\lambda^{\vee}_{\pm}\sqcup\lambda^{\vee}_{\text{uni}}\text{ and }{}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\mu_{\pm}+\mu_{\text{uni}}.$
Then
$d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ $=d^{\vee}(\lambda^{\vee}_{\pm})+(\lambda^{\vee}_{\text{uni}})^{t}\geq\mu_{\pm}+\mu_{\text{uni}},$
since by Propositions 4.5 and 4.4, we know that
$d^{\vee}(\lambda^{\vee}_{\pm})\geq\mu_{\pm}\text{ and }(\lambda^{\vee}_{\text{uni}})^{t}\geq\mu_{\text{uni}}.$
In the case when $n_{2}=0$ (hyperspecial), equality holds for both, therefore for the combination. ∎
4.3. Unramified unitary groups
In this case, the calculation is simply that from Proposition 4.4, so there is nothing more to check.
4.4. Special orthogonal groups
The calculation is similar to that for the symplectic group. The part of the calculation pertaining to the polynomials $P$ of degree strictly greater than $1$ is identical to Proposition 4.4. Thus, we only need to prove the analogue of Proposition 4.5 for orthogonal groups.
Proposition 4.7.
Let $\pi$ be a depth-zero supercuspidal $SO(m,\mathsf{k})$-representation. Suppose there are no endoscopic unitary factors, i.e., no polynomials $P$ with $\deg P>1$ appear. Then,
  1. (1)
    the nilpotent Langlands parameter ${\mathbb{O}}^{\vee}_{\pi}$ is given by the strings (2o) and (3o) in section 4.1.
  2. (2)
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\mathsf{WF}_{1}\sqcup\mathsf{WF}_{2},$
    where
    $\mathsf{WF}_{i}=\begin{cases}[2(m^{(i)}_{+}+m^{(i)}_{-})+1]\sqcup[2|m^{(i)}_{+}-m^{(i)}_{-}|-1],&n_{i}^{an}=1,\\ [2(m^{(i)}_{+}+m^{(i)}_{-})-1]\sqcup[2|m^{(i)}_{+}-m^{(i)}_{-}|-1]&n_{i}^{an}=0,2.\end{cases}$
  3. (3)
    Conjecture 1.6 holds in this case.
Proof.
Claim (1) is only recorded here for convenience. For (2), we need to compute the Kawanaka wavefront set of the cuspidal representations on each factor of ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}$. Consider one factor $SO(n+n^{an},n,{\mathbb{F}}_{q})$. We only need to compute the orthogonal nilpotent orbit that is attached via Springer’s correspondence to the $j$-induced representation
$\sigma=\begin{cases}j_{W(B)\times W(B)}^{W(B)}\left(((1,2,\dots,m_{+})\times(1,2,\dots,m_{+}))\boxtimes(1,2,\dots,m_{-})\times(1,2,\dots,m_{-}))\right),&n^{an}=1,\\ j_{W(D)\times W(D)}^{W(D)}\left(((1,2,\dots,m_{+})\times(1,2,\dots,m_{+}-1))\boxtimes(1,2,\dots,m_{-})\times(1,2,\dots,m_{-}-1))\right),&n^{an}=0,2.\end{cases}$
As in the symplectic case, by the formulae in [Lu3], this equals
$\sigma=\begin{cases}((1,2,\dots,m_{+})+(1,2,\dots,m_{-}))\times((1,2,\dots,m_{+})+(1,2,\dots,m_{-})),&n^{an}=1,\\ ((1,2,\dots,m_{+})+(1,2,\dots,m_{-}))\times((1,2,\dots,m_{+}-1)+(1,2,\dots,m_{-}-1)),&n^{an}=0,2.\end{cases}$
Applying the combinatorial algorithm for the Springer representation for type $B$, we find that the corresponding orthogonal nilpotent orbit is
$\begin{cases}[2(m_{+}+m_{-})+1]\sqcup[2|m_{+}-m_{-}|-1],&n^{an}=1,\\ [2(m_{+}+m_{-})-1]\sqcup[2|m_{+}-m_{-}|-1],&n^{an}=0,2.\end{cases}$
Then (2) follows. Claim (3) is verified in the same way as for the symplectic case.
By combining Proposition 4.7 with Proposition 4.4, it follows, just as in the proof of Theorem 4.6 that Conjecture 1.6 holds for all depth-zero supercuspidals representations of a special orthogonal group.
4.5. Ramified unitary groups
For the ramified unitary groups, the finite group ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})$ on which the cuspidal representation lives is a product of a symplectic and special orthogonal group. For notational purposes, assume ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}({\mathbb{F}}_{q})=Sp(2n_{1},{\mathbb{F}}_{q})\times SO(n_{2}+n_{2}^{an},n_{2},{\mathbb{F}}_{q})$. We only record the analogue of Propositions 4.5 and 4.7. The proof is simply a combination of the symplectic and orthogonal cases discussed before.
Proposition 4.8.
Let $\pi$ be a depth-zero supercuspidal representation of a ramified unitary group induced from ${{\mathcal{G}}}^{\circ}_{n_{1},n_{2}}$ as above. Suppose there are no endoscopic unitary factors, i.e., no polynomials $P$ with $\deg P>1$ appear. Then,
  1. (1)
    the nilpotent Langlands parameter ${\mathbb{O}}^{\vee}_{\pi}$ is given by the strings (2r) and (3r) in section 4.1.
  2. (2)
    ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)=\mathsf{WF}_{1}\sqcup\mathsf{WF}_{2},$
    where
    $\mathsf{WF}_{1}=[2(m_{+}^{(1)}+m_{-}^{(1)})]\sqcup\begin{cases}[2(m_{+}^{(1)}-m_{-}^{(1)})],&m_{+}^{(1)}\geq m_{-}^{(1)},\\ [2(m_{-}^{(1)}-m_{+}^{(1)}-1)],&m_{+}^{(1)}<m_{-}^{(1)}\end{cases}$
    and
    $\mathsf{WF}_{2}=\begin{cases}[2(m^{(2)}_{+}+m^{(2)}_{-})+1]\sqcup[2|m^{(2)}_{+}-m^{(2)}_{-}|-1],&n_{2}^{an}=1,\\ [2(m^{(2)}_{+}+m^{(2)}_{-})-1]\sqcup[2|m^{(2)}_{+}-m^{(2)}_{-}|-1]&n_{2}^{an}=0,2.\end{cases}$
  3. (3)
    Conjecture 1.6 holds in this case.
5. Tempered L-packets of $\mathbf{G_{2}}$
Langlands correspondences for $G_{2}$ were recently constructed in [AX] and [GS1, GS2]. We verify Conjecture 1.6 for $G_{2}$ using the explicit description of L-packets in [AX].
5.1. The verification of the main conjectures for $G_{2}$
Using the explicit correspondence and the results on the wavefront sets of $G_{2}$ from [JLS] and [LoSa], we can verify our main conjectures. Recall that $G_{2}$ has five geometric orbits labelled in the Bala-Carter classification, as follows (in the closure order):
${\mathrm{G}_{2}},\quad{\mathrm{G}_{2}(a_{1})},\quad{\widetilde{A}_{1}},\quad{A_{1}},\quad{1}.$
Lemma 5.1.
If $\pi$ is not trivial and not generic, then ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)={\mathrm{G}_{2}(a_{1})}.$
Proof.
By [JLS, §12 and Table $G_{2}$, p. 437], the only possible geometric wavefront sets in $G_{2}$ are ${\mathrm{G}_{2}}$, ${\mathrm{G}_{2}(a_{1})}$, ${A_{1}}$, and 11{1}1. By [LoSa, Theorem 1.2], the minimal orbit ${A_{1}}$ does not occur as a wavefront set. Hence ${\mathrm{G}_{2}}$, ${\mathrm{G}_{2}(a_{1})}$, and 11{1}1 are the only possibilities. The wavefront set of $\pi$ is the regular orbit ${\mathrm{G}_{2}}$ if and only if $\pi$ is generic. On the other hand, the wavefront set is $1$ if and only if $\pi$ is an unramified representation, hence trivial. ∎
Theorem 5.2.
Let $\pi$ be a tempered irreducible representation of $G_{2}$. Let ${\mathbb{O}}_{\pi}^{\vee}\in\{{\mathrm{G}_{2}},{\mathrm{G}_{2}(a_{1})},{\widetilde{A}_{1}},{A_{1}},{1}\}$ be the nilpotent part of the Langlands parameter of $\pi$. If ${\mathbb{O}}_{\pi}^{\vee}$ is not the trivial orbit, ${\mathsf{AZ}}(\pi)$ is not generic and $\mathsf{WF}({\mathsf{AZ}}(\pi))={\mathrm{G}_{2}(a_{1})}.$
Proof.
We fix a maximal ${\mathsf{k}}$-split torus $T$, a Borel subgroup $B$ with $T\subset B$. Let $\Delta=\{\alpha,\beta\}$ be the set of simple roots associated to $B$ where $\alpha$ is a short root. Let $P_{\gamma}=M_{\gamma}U_{\gamma}$, $\gamma\in\Delta\cup\{\emptyset\}$ be the parabolic subgroup associated to $\gamma$. We identify $T$ with ${\mathsf{k}}^{\times}\times{\mathsf{k}}^{\times}$ via $M_{\alpha}\simeq GL_{2}$. We also adopt the notation from [GS1]:
  • $I(s_{1},s_{2},\chi_{1},\chi_{2})=i_{B}^{G}(\nu_{\mathsf{k}}^{s_{1}}\chi_{1}\otimes\nu_{\mathsf{k}}^{s_{2}}\chi_{2})$ where $\chi_{i}$ is a character of ${\mathsf{k}}^{\times}$ and $\nu_{\mathsf{k}}=|\cdot|_{\mathsf{k}}.$
  • $I_{P_{\gamma}}(s,\tau)=i_{P_{\gamma}}^{G}(\nu^{s}\otimes\tau)$ where $\nu=\nu_{\mathsf{k}}\circ\det$ and $\tau$ is a representation of $M_{\gamma}\simeq GL_{2}$.
  • $J_{R}(\cdots)$ where $R$ is a parabolic denotes a Langlands quotient of $I_{R}(\cdots)$.
  • $G:=G_{2}$, $P:=P_{\alpha}$, $Q:=P_{\beta}$
Thanks to the classification of unitary dual of $G_{2}$ by Muic ([Mu]), we can make an explicit calculation of the ${\mathsf{AZ}}$ duals. We also need the following facts:
  • (i)
    The ${\mathsf{AZ}}$ dual of an irreducible representation $\pi$ is irreducible.
  • (ii)
    The ${\mathsf{AZ}}$ dual map is compatible with parabolic inductions, that is, $i_{P}^{G}{\mathsf{AZ}}(\tau)={\mathsf{AZ}}(i_{P}^{G}\tau)$, where $\tau$ is an admissible representation of $M$.
  • (iii)
    If $\pi$ is irreducible, $\pi$ and ${\mathsf{AZ}}(\pi)$ have the same supercuspidal support. Hence, if $\tau$ is a supercuspidal representation of $M_{\gamma}$, ${\mathsf{AZ}}(i_{P_{\gamma}}^{G}\tau)=i_{P_{\gamma}}^{G}\tau$ in the Grothendieck group $K(G_{2})$ of smooth representations.
  • (iv)
    If $\pi$ is a nontrivial non-generic irreducible representation of $G_{2}$, $\mathsf{WF}(\pi)$ is ${\mathrm{G}_{2}(a_{1})}$.
We now explain the calculation of the ${\mathsf{AZ}}$ map. Since subquotients of the parabolically induced presentations are preserved by the ${\mathsf{AZ}}$ map, we can consider the ${\mathsf{AZ}}$ map restricted to a finite set of representations with the same cuspidal support. We will show two cases since the rest of calculation is similar: In the following, $=$, $+$ are operations in $K(G_{2})$.
Case 1. Subquotients of $I(1,0,1_{{\mathsf{k}}^{\times}},1_{{\mathsf{k}}^{\times}})$: By [GS1, Prop. 3.1-3.2], we have in $K(G_{2})$
$\begin{array}[]{lcl}I(1,0,1_{{\mathsf{k}}^{\times}},1_{{\mathsf{k}}^{\times}})&=&I_{P}(\frac{1}{2},\delta(1))+I_{P}(\frac{1}{2},1_{GL_{2}})\\ &=&I_{Q}(\frac{1}{2},\delta(1))+I_{Q}(\frac{1}{2},1_{GL_{2}})\\ I_{P}(\frac{1}{2},\delta(1))&=&\pi_{\textup{gen}}(1)+J_{P}(\frac{1}{2},\delta(1))+J_{Q}(\frac{1}{2},\delta(1))\\ I_{P}(\frac{1}{2},1_{GL_{2}})&=&\pi_{\textup{deg}}(1)+J_{Q}(1,\pi(1,1))+J_{Q}(\frac{1}{2},\delta(1))\\ I_{Q}(\frac{1}{2},\delta(1))&=&\pi_{\textup{gen}}(1)+\pi_{\textup{deg}}(1)+J_{Q}(\frac{1}{2},\delta(1))\\ I_{Q}(\frac{1}{2},1_{GL_{2}})&=&J_{Q}(\frac{1}{2},\delta(1))+J_{Q}(1,\pi(1,1))+J_{P}(\frac{1}{2},\delta(1))\end{array}$
where $\delta(1)$ is the Steinberg representation of $GL_{2}$. Since ${\mathsf{AZ}}_{GL_{2}}(\delta(1))=1_{GL_{2}}$, we have
$\begin{array}[]{lccl}{\mathsf{AZ}}:&\left\{\pi_{\textup{gen}}(1),\ J_{P}(\frac{1}{2},\delta(1)),\ J_{Q}(\frac{1}{2},\delta(1))\right\}&\Leftrightarrow&\left\{\pi_{\textup{deg}}(1),\ J_{Q}(1,\pi(1,1)),\ J_{Q}(\frac{1}{2},\delta(1))\right\}\\ &\left\{\pi_{\textup{gen}}(1),\ \pi_{\textup{deg}}(1),\ J_{Q}(\frac{1}{2},\delta(1))\right\}&\Leftrightarrow&\left\{J_{Q}(\frac{1}{2},\delta(1)),\ J_{Q}(1,\pi(1,1)),\ J_{P}(\frac{1}{2},\delta(1))\right\}\end{array}$
Then, the only possible matching is
$\pi_{\textup{gen}}(1)\overset{{\mathsf{AZ}}}{\leftrightarrow}J_{Q}(1,\pi(1,1)),\quad\pi_{\textup{deg}}(1)\overset{{\mathsf{AZ}}}{\leftrightarrow}J_{P}(1/2,\delta(1)),\quad J_{Q}(1/2,\delta(1))\overset{{\mathsf{AZ}}}{\leftrightarrow}J_{Q}(1/2,\delta(1))$
Case 2. Subquotients of $I(1,0,\chi,\chi)=I(1,0,\chi^{-1},\chi^{-1})$ with $\chi^{3}=1$: The above method gives only
${\mathsf{AZ}}:\{\pi_{\textup{gen}}(\chi)=\pi_{\textup{gen}}(\chi^{-1}),J_{P}(1/2,\delta(\chi)\}\ \Leftrightarrow\ \{J_{P}(1/2,\delta(\chi^{-1}),J_{Q}(1/2,\pi(\chi,\chi^{-1}))\}$
However, using $I(1,0,\chi,\chi)=I(1,0,\chi^{-1},\chi^{-1})$, we can nail it down to
$\pi_{\textup{gen}}(\chi)\overset{{\mathsf{AZ}}}{\leftrightarrow}J_{Q}(1/2,\pi(\chi,\chi^{-1})),\qquad J_{P}(1/2,\delta(\chi)\overset{{\mathsf{AZ}}}{\leftrightarrow}J_{P}(1/2,\delta(\chi^{-1}))$
$\pi$ ${\mathbb{O}}_{\pi}$ $d^{\vee}({\mathbb{O}}_{\pi})$ ${\mathsf{AZ}}(\pi)$ $\mathsf{WF}({\mathsf{AZ}}(\pi))$ reference
$I(2,1,1_{{\mathsf{k}}^{\times}},1_{{\mathsf{k}}^{\times}})$ $St_{G_{2}}$ ${\mathrm{G}_{2}}$ 11{1}1 $1_{G_{2}}$ 11{1}1 Prop. 3.1 (ii)-(iii) Prop. 3.2 (ii)-(iii) [GS1]
$J_{Q}(\frac{5}{2},\delta(1))$ ${\widetilde{A}_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{P}(\frac{3}{2},\delta(1))$ ${\mathrm{G}_{2}(a_{1})}$
$J_{P}(\frac{3}{2},\delta(1))$ ${A_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(\frac{5}{2},\delta(1))$ ${\mathrm{G}_{2}(a_{1})}$
$1_{G_{2}}$ 11{1}1 ${\mathrm{G}_{2}}$ $St_{G_{2}}$ ${\mathrm{G}_{2}}$
$I(1,0,1_{{\mathsf{k}}^{\times}},1_{{\mathsf{k}}^{\times}})$ $\pi_{\textup{gen}}(1)$ ${\mathrm{G}_{2}(a_{1})}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(1,\pi(1,1)$ ${\mathrm{G}_{2}(a_{1})}$
$\pi_{\textup{deg}}^{r}(1)$ ${\mathrm{G}_{2}(a_{1})}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{P}(\frac{1}{2},\delta(1))$ ${\mathrm{G}_{2}(a_{1})}$
$J_{Q}(\frac{1}{2},\delta(1))$ ${\widetilde{A}_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(\frac{1}{2},\delta(1))$ ${\mathrm{G}_{2}(a_{1})}$
$J_{P}(\frac{1}{2},\delta(1))$ ${A_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $\pi_{\textup{deg}}^{r}(1)$ ${\mathrm{G}_{2}(a_{1})}$
$J_{Q}(1,\pi(1,1))$ 11{1}1 ${\mathrm{G}_{2}}$ $\pi_{\textup{gen}}(1)$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I(1,0,\chi\otimes\chi)\\ \chi^{2}=1\end{array}$ $\pi_{\textup{gen}}(\chi)$ ${\mathrm{G}_{2}(a_{1})}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(\frac{1}{2},\pi(1,\chi))$ ${\mathrm{G}_{2}(a_{1})}$
$J_{Q}(\frac{1}{2},\delta(\chi)$ ${\widetilde{A}_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{P}(\frac{1}{2},\delta(\chi))$ ${\mathrm{G}_{2}(a_{1})}$
$J_{P}(\frac{1}{2},\delta(\chi)$ ${A_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(\frac{1}{2},\delta(\chi))$ ${\mathrm{G}_{2}(a_{1})}$
$J_{Q}(\frac{1}{2},\pi(1,\chi))$ 11{1}1 ${\mathrm{G}_{2}}$ $\pi_{\textup{gen}}(\chi)$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I(1,0,\chi\otimes\chi)\\ \chi^{3}=1\end{array}$ $\pi_{\textup{gen}}(\chi)$ ${\mathrm{G}_{2}(a_{1})}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(\frac{1}{2},\pi(1,\chi))$ ${\mathrm{G}_{2}(a_{1})}$
$J_{P}(\frac{1}{2},\delta(\chi)$ ${A_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{P}(\frac{1}{2},\delta(\chi^{-1}))$ ${\mathrm{G}_{2}(a_{1})}$
$J_{P}(\frac{1}{2},\delta(\chi^{-1})$ ${A_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(\frac{1}{2},\delta(\chi))$ ${\mathrm{G}_{2}(a_{1})}$
$J_{Q}(1,\pi(\chi,\chi^{-1}))$ 11{1}1 ${\mathrm{G}_{2}}$ $\pi_{\textup{gen}}(\chi)$ ${\mathrm{G}_{2}}$
$i_{B}^{G}\chi$ irred. $i_{B}^{G}\chi$ 11{1}1 ${\mathrm{G}_{2}}$ $i_{B}^{G}\chi$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I_{P}(\frac{1}{2},\tau)\\ \tau\simeq\tau^{\vee},\ \omega_{\tau}=1\end{array}$ $\delta_{P}(\tau)$ ${A_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{P}(\frac{1}{2},\tau)$ ${\mathrm{G}_{2}(a_{1})}$ [GS1, Prop. 3.1-(i)]
$J_{P}(\frac{1}{2},\tau)$ 11{1}1 ${\mathrm{G}_{2}}$ $\delta_{P}(\tau)$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I_{P}(0,\tau),\tau\simeq\tau^{\vee}\\ \omega_{\tau}\neq 1\end{array}$ $I_{P}(\tau)_{\textup{gen}}$ 11{1}1 ${\mathrm{G}_{2}}$ $I_{P}(\tau)_{\textup{deg}}$ ${\mathrm{G}_{2}(a_{1})}$ [GS1, Prop. 3.1-(ii)]
$I_{P}(\tau)_{\textup{deg}}$ 11{1}1 ${\mathrm{G}_{2}}$ $I_{P}(\tau)_{\textup{gen}}$ ${\mathrm{G}_{2}}$
$i_{P}^{G}\tau$ irred. $i_{Q}^{G}\tau$ 11{1}1 ${\mathrm{G}_{2}}$ $i_{P}^{G}\tau$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I_{Q}(\frac{1}{2},\tau)\\ \tau\simeq\tau^{\vee},\ \omega_{\tau}=1\end{array}$ $\delta_{Q}(\tau)$ ${\widetilde{A}_{1}}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(\frac{1}{2},\tau)$ ${\mathrm{G}_{2}(a_{1})}$ [GS1, Prop. 3.2-(i)]
$J_{Q}(\frac{1}{2},\tau)$ 11{1}1 ${\mathrm{G}_{2}}$ $\delta_{Q}(\tau)$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I_{Q}(1,\tau),\ \tau\simeq\tau^{\vee}\\ \tau\textrm{ dihedral},\ \omega_{\tau}\neq 1\end{array}$ $\pi_{\textup{gen}}(\tau)$ ${\mathrm{G}_{2}(a_{1})}$ ${\mathrm{G}_{2}(a_{1})}$ $J_{Q}(1,\tau)$ ${\mathrm{G}_{2}(a_{1})}$ [GS1, Prop. 3.2-(ii)]
$J_{Q}(1,\tau)$ 11{1}1 ${\mathrm{G}_{2}}$ $\pi_{\textup{gen}}(\tau)$ ${\mathrm{G}_{2}}$
$\begin{array}[]{c}I_{Q}(0,\tau),\tau\simeq\tau^{\vee}\\ \omega_{\tau}\neq 1\end{array}$ $I_{Q}(\tau)_{\textup{gen}}$ 11{1}1 ${\mathrm{G}_{2}}$ $I_{Q}(\tau)_{\textup{deg}}$ ${\mathrm{G}_{2}(a_{1})}$ [GS1, Prop. 3.2-(iii)]
$I_{Q}(\tau)_{\textup{deg}}$ 11{1}1 ${\mathrm{G}_{2}}$ $I_{Q}(\tau)_{\textup{gen}}$ ${\mathrm{G}_{2}}$
$i_{Q}^{G}\tau$ irred. $i_{Q}^{G}\tau$ 11{1}1 ${\mathrm{G}_{2}}$ $i_{Q}^{G}\tau$ ${\mathrm{G}_{2}}$
$\pi$, nonsingular s.c. $\pi$ 11{1}1 ${\mathrm{G}_{2}}$ $\pi$ ${\mathrm{G}_{2}(a_{1})}$ [Kal]
or ${\mathrm{G}_{2}}$
$\pi$, singular s.c. $\pi$ $\begin{array}[]{c}{\widetilde{A}_{1}},\ {A_{1}}\\ \textrm{or }{\mathrm{G}_{2}(a_{1})}\end{array}$ ${\mathrm{G}_{2}(a_{1})}$ $\pi$ ${\mathrm{G}_{2}(a_{1})}$
Table 1. Irrreducible tempered representations of $G_{2}$ and their ${\mathsf{AZ}}$-duals
Corollary 5.3.
Conjecture 1.9 holds for any irreducible tempered representation $\pi$ on $G_{2}$.
Proof.
This is immediate from the previous results. ∎
5.2. Depth-zero representations
In the case of the depth-zero representations, it is instructive to compute the unramified wavefront set as well in order to test several possible refinements of our main conjecture.
Proposition 5.4.
The canonical unramified wavefront sets of the depth-zero supercuspidal representations in the parametrization of [AX] are given in Table 2. In particular, Conjecture 1.1 holds for the depth-zero supercuspidals of $G_{2}$.
Proof.
The depth-zero supercuspidals are parameterized in loc. cit. as in the following Table 2. From the realization of the representations by compact induction, we can immediately compute the canonical unramified ${}^{K}\underline{\mathsf{WF}}(\pi)$ (in the sense of Okada [Ok]) and the geometric wavefront sets, via Theorem 3.1. Then one verifies by inspection the conjecture. The unipotent supercuspidal representations have already been treated in [CMBO2]. We recall that
$d^{\vee}(G_{2})=0,\quad d^{\vee}(0)=G_{2},\text{ and }d^{\vee}({\mathbb{O}}^{\vee})=G_{2}(a_{1}),$
if ${\mathbb{O}}^{\vee}=G_{2}(a_{1})$, $\widetilde{A}_{1}$, or $A_{1}$.
We exemplify the calculation of the wavefront set for $\pi$ in the sixth row, for which ${\mathbb{O}}^{\vee}=A_{1}$. Here ${\mathfrak{g}^{\vee}}^{\varphi(I_{\mathsf{k}})}=sl(3,{\mathbb{C}})$. The supercuspidal representation is compactly induced from the hyperspecial parahoric from the inflation of a cuspidal representation $\mu$ of $G_{2}({\mathbb{F}}_{q})$. The cuspidal representation $\mu$ belongs to the Lusztig series corresponding to a semisimple element $\tau\in G_{2}^{\vee}({\mathbb{F}}_{q})$, such that the centralizer of $\tau$ is $SU(3)$. (Recall that we need $q\equiv-1(3).$) In $SU(3)$, the unipotent class parametrizing $\tau$ is $(12)$, so to find the Kawanaka wavefront set of $\mu$ we need the $j$-induced
$j_{S_{3}}^{W(G_{2})}((12))=2_{1},$
the reflection representation of $W(G_{2})$. The corresponding nilpotent orbit via Springer’s correspondence is $G_{2}(a_{1}).$ This is the (geometric) Kawanaka wavefront set of $\mu$. Since the parahoric is the hyperspecial, it follows that the canonical unramified wavefront set of $\pi$ is indeed $(G_{2}(a_{1}),1)$.
${\mathbb{O}}^{\vee}$ $\varphi(I_{\mathsf{k}})$ $\varphi(\langle\mathsf{Fr}\rangle)$ $G_{x,0}/G_{x,0^{+}}$ $\pi$ ${}^{K}\underline{\mathsf{WF}}(\pi)$ $d^{\vee}({\mathbb{O}}^{\vee})$ $d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))$
$G_{2}(a_{1})$ $1$ $1$ $G_{2}$ $G_{2}(1)$ $(G_{2}(a_{1}),1)$ $G_{2}(a_{1})$ $G_{2}(a_{1})$
$\mu_{2}$ $G_{2}(-1)$
$\mu_{3}$ $G_{2}(\omega),G_{2}(\omega^{2})$
$G_{2}(a_{1})$ $\mu_{2}$ $SO_{4}$ singular $(G_{2}(a_{1}),(12))$ $G_{2}(a_{1})$ $\widetilde{A}_{1}$
$G_{2}(a_{1})$ $\mu_{3}$ $SL_{3}$ singular, $q\equiv 1(3)$ $(G_{2}(a_{1}),(123))$ $G_{2}(a_{1})$ $A_{1}$
$A_{1}$ $\mu_{3}$ $G_{2}$ singular from $SU(3)$ $(G_{2}(a_{1}),1)$ $G_{2}(a_{1})$ $G_{2}(a_{1})$
$q\equiv-1(3)$
$1$ $G_{2}$ nonsingular $(G_{2},1)$ $G_{2}$ $1$
$SL_{3}$ toral $(G_{2}(a_{1}),(123))$ $G_{2}$ $A_{1}$
$SO_{4}$ $(G_{2}(a_{1}),(12))$ $G_{2}$ $\widetilde{A}_{1}$
Table 2. Depth-zero supercuspidal representations of $G_{2}$
To account for an example of an anti-tempered unipotent $E_{7}$-representation for which ${}^{\bar{\mathsf{k}}}\mathsf{WF}(\pi)<d^{\vee}({\mathbb{O}}^{\vee}_{\pi})$ (strict inequality), Okada proposed another, more precise form of the relation between the wavefront set of unipotent representations and the Langlands parameter:
$d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{FT}}.$ (5.1)
This holds for unipotent representations with real infinitesimal character [CMBO1], and there are no known counterexamples for unipotent representations. However, for more general representations, this equality can’t always hold as stated, for example, because of the case of nongeneric regular depth-zero supercuspidals, see section 3.1.
A natural attempt to fix it is to intersect with ${\mathfrak{h}}^{\vee}={\mathfrak{g}^{\vee}}^{\varphi(I_{\mathsf{k}})}$, i.e., to expect that $\overline{d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))}\cap{\mathfrak{h}}^{\vee}$ is an irreducible $H^{\vee}(s)$-variety, $s=\varphi(\mathsf{Fr})$, hence the closure of a nilpotent $H^{\vee}(s)$-orbit ${{\mathcal{O}}}^{\vee}$ such that
${\mathbb{O}}^{\vee}=G^{\vee}\cdot{{\mathcal{O}}}^{\vee}_{\mathsf{FT}}.$ (5.2)
Often this seems to be the case (for example, this resolves trivially the case of regular depth-zero supercuspidals), but again not always, as the $G_{2}$ examples above show.
More precisely, consider $\pi$ in the sixth row with ${\mathbb{O}}^{\vee}_{\pi}=A_{1}$ and ${}^{K}\underline{\mathsf{WF}}(\pi)=(G_{2}(a_{1}),1)$ and $d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))=G_{2}(a_{1})$. The intersection $\overline{d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))}\cap{\mathfrak{h}}^{\vee}$ is the principal orbit in $sl(3,{\mathbb{C}})$. However, the Langlands nilpotent orbit ${\mathbb{O}}^{\vee}=A_{1}$ comes from the subregular orbit in $sl(3,{\mathbb{C}})$.
Interestingly, the $G_{2}$ depth-zero supercuspidals in Table 2 seem to indicate that there is no generalization of (5.1) even if one would replace the equality with an inequality, i.e.:
$d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))<{\mathbb{O}}^{\vee},\quad$ $\pi\text{ in the fourth or fifth rows},$
$d_{S}({}^{K}\underline{\mathsf{WF}}(\pi))>{\mathbb{O}}^{\vee},\quad$ $\pi\text{ in the sixth row}.$
This calculation allows for an alternative verification of the conjecture in the case of depth-zero L-packets that contain a supercuspidal representation.
Proposition 5.5.
Conjecture 1.9 holds for all depth-zero discrete series L-packets of $G_{2}$ that contain a supercuspidal representation in the parametrization of [AX].
Proof.
The parametrization of the unipotent discrete series L-packets agrees with the Kazhdan-Lusztig parametrization [KL], hence the claim is already covered by [CMBO1] in the case of real infinitesimal character.
For the regular supercuspidal L-packets, the claim is immediate and has already been discussed in section 3.1
For the non-unipotent mixed L-packets, notice that Table 2 shows that the conjectural upper bound is attained by the supercuspidal in the L-packet. So it only remains to check that if $\pi$ is a non-supercuspidal representation in the L-packet, then ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))\leq d^{\vee}({\mathbb{O}}^{\vee})$. In all cases, $d^{\vee}({\mathbb{O}}^{\vee})=G_{2}(a_{1}).$ So the only way that the conjecture could fail is if ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=G_{2}$, which would make ${\mathsf{AZ}}(\pi)$ generic. But from the classification, we know that $\pi$ itself is generic, in fact $\pi$ is the generic composition factor of a generalized principal series representation that has two composition factors: $\pi$ and $\pi^{\prime}$. Then clearly ${\mathsf{AZ}}(\pi)=\pi^{\prime}$ which is not generic. This concludes the proof.
In fact, the previous proof can be adapted to conclude the following result (which is already known in more generality from [JLS, LoSa]).
Proposition 5.6.
Let $\pi$ be an irreducible unipotent representation of $G_{2}$. Then either $\pi$ is generic, or $\pi$ is trivial, or ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))=G_{2}(a_{1})$.
Proof.
Suppose $\pi$ is not generic. Then ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))\leq G_{2}(a_{1})$. Consider the parahoric restriction $\mathsf{res}_{K_{0}}(\pi)$ of $\pi$ with respect to the maximal hyperspecial subgroup $K_{0}$. This is a unipotent representation of $G_{2}(\mathbb{F}_{q})$ and as such, its geometric Kawanaka wavefront set $\mathsf{WF}_{q}(\mathsf{res}_{K_{0}}(\pi))$ is special. The only special orbits in $G_{2}(\overline{\mathbb{F}}_{q})$ are the regular, subregular, and the trivial orbit. The only way $\mathsf{WF}_{q}(\mathsf{res}_{K_{0}}(\pi))$ is the trivial orbit is if $\mathsf{res}_{K_{0}}(\pi)$ is a trivial $G_{2}(\mathbb{F}_{q})$-representation. In particular, this means that $\pi$ is a $K_{0}$-spherical representation and in the equivalence of categories with the Iwahori-Hecke algebra, the corresponding Iwahori-Hecke algebra module only contains the trivial $K_{0}$-type. This implies that $\pi$ must be an unramified character, but since the group is $G_{2}$, it must be the trivial representation.
By [Ok, Theorem 5.4], ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))$ is greater than the lift to $G$ of $\mathsf{WF}_{q}(\mathsf{res}_{K_{0}}(\pi))$. So if $\pi$ is not trivial, it follows that ${}^{\bar{\mathsf{k}}}\mathsf{WF}({\mathsf{AZ}}(\pi))\geq G_{2}(a_{1})$ and this completes the proof.